OpenStudy (eric_d):

Express x^2+1/(x-1)(x+1)^3 in partial fraction(EXAMPLE)

OpenStudy (eric_d):

Solution: x^2+1=A(x+1)^3+b(x-1)(x+1)^2+c(x-1)(x+1)+D(x-1) When x=1, 2=8A A=1/4 When x=-1,2=-2D D=-1

OpenStudy (eric_d):

@ganeshie8

OpenStudy (eric_d):